# Spiking Neural Network Architecture Search: A Survey

### v1

@@ III-A Unique Challenges of Applying NAS to SNNs
First, while standard NAS relies on gradient-based optimization, SNNs’ discrete spike functions are inherently non-differentiable. Though surrogate gradient methods exist, they introduce approximation errors and computational overhead that impact large-scale architecture searches [74]. The prevalence of spike-based learning rules like STDP, distinct from gradient-based approaches, further complicates traditional NAS application.
@@
Third, neuromorphic hardware considerations introduce unique constraints such as limited precision and asynchronous computation [75]. Traditional NAS methods, which do not account for these constraints, may produce architectures incompatible with neuromorphic platforms.
@@
III-B Search Space
The search space for a NAS method is the collection of feasible network configurations, typically defined by parameters such as the number and type of layers, and the layer hyperparameters (e.g., kernel size, filters, or number of neurons). Because these dimensions combine to yield billions of potential architectures, the search space can be prohibitively large and often becomes the primary bottleneck for architecture search performance [76]. Moreover, different architectural choices can significantly impact hardware accelerators, affecting power efficiency, memory usage, and throughput [60, 37]. Consequently, it is important to design the search space to account for practical hardware constraints. Balancing performance and hardware friendliness ensures that the discovered architectures are accurate and can be efficiently implemented in hardware.
The search can be simplified, and the search space reduced by incorporating prior knowledge about well-performing architectural patterns. However, doing so may introduce human bias that undermines the discovery of novel solutions. On the other hand, overly constraining the search space might exclude high-performing architectures, resulting in suboptimal performance. Hence, selecting an appropriate search space remains a critical challenge in SNNaS, requiring consideration of both algorithmic and hardware factors.
@@ III-B Search Space
Global Search Space
@@
Global search spaces consider the entire network configuration at once. They allow for tremendous diversity but can be extremely large. For instance, MixedSNN [19] searches among different spiking neuron models and thresholds, and Auto-Spikformer [17] combines Transformer and SNN parameters. Methods such as Zhou et al.[13], Liu and Yi[20], Yan et al.[24], AutoST[28], GAQ-SNN [16], and ANAS [77] illustrate the complexity of global search—spanning from trillions of spiking configurations to hardware-related parameters like processing elements and buffer sizes.
Sequential Search Space
Inspired by early NAS works [78, 79], sequential designs treat a network as a layer-by-layer sequence of operations (e.g., convolution parameters). While powerful, these methods can require enormous compute resources. Sequential spaces remain largely unexplored in SNNaS..
Cell-Based Search Space
@@
Cell-based approaches define networks at macro and micro levels. Macro-architectures arrange repeated cells, while micro-architectures define operations within each cell. SNASNet [8], based on NAS-Bench-201 [80], pioneered cell-based SNN search using directed acyclic graphs (DAGs) of operations (Zeroize, SkipCon, $\times1$Conv, $\times3$Conv, $\times3$AvgPool). Extensions include SSTNAS [81], SpikeNAS [12], larger-kernel exploration [21], hardware-friendly constraints [22], and variations in cell design [14]. SpikeDHS [15, 82] also employs DAG-based cells. An example cell-based diagram is shown in Figure 3.
A diagram titled ‘Cell’ that illustrates a cell-based search space. Inside a rounded rectangle are four black nodes arranged left-to-right (label them A, B, C, D). The cell is a directed acyclic graph, and the directed colored arrows between the nodes represent operations. A legend on the right maps arrow colors to operations: blue represents zeroize, yellow represents skip-connect, red represents one-by-one convolution, green represents three-by-three convolution, and purple represents three-by-three average pooling. The node connections shown are: A to B with a yellow arrow (skip-connect), A to C with a blue arrow (zeroize), B to C with a green arrow (three-by-three convolution), C to D with a purple arrow (three-by-three average pooling), A to D with a red curved arrow (one-by-one convolution), and B to D with a blue arrow (another zeroize operation).
@@
Figure 3: Example of a cell with four nodes and five possible operations. The cell is a directed acyclic graph, with each edge representing an operation [80].
Hierarchical Structure Search Space
@@
Hierarchical NAS arranges architectures in multi-level motifs, mirroring modular designs like VGGNet [83] or ResNet [11]. MSE-NAS [23] organizes SNNs into microscopic, mesoscopic, and macroscopic levels; AutoSNN [7] and ANCoEF [84] combine macro-level backbones with micro-level spiking blocks. SpikeDHS [15, 82] extends from cell-level to layer-level search, and LitE-SNN [25] adds pruning and mixed-precision quantization. SpikeExplorer [26] employs Bayesian optimization for multi-objective search. DESPINE [27] and NeuEvo [18] further emphasize biologically inspired motifs, excitatory/inhibitory neurons, and feedback connections. Figure 4 illustrates a typical three-level hierarchy.
[…]
@@
Memory Bank Representation Search Space
Instead of conventional graphs, SMASH [86] encodes a network as a memory bank system where each operation reads/writes from distinct memory blocks. Figure 5 contrasts single-branch and branching designs. Alhough SMASH’s representation reduces training overhead, it remains unexplored in SNNaS.
Diagram comparing memory-bank usage in single-branch and branching networks. Color/shape key: gray rectangles are labeled as memory banks; tall yellow rectangles labeled ‘conv’ represent convolution operations; blue arrows represent reads from a bank into a convolution; red arrows represent writes from a convolution back to a bank. A single-branch network with one memory bank is illustrated on the left. A convolution operation reads from this bank and writes its output back into the same bank, indicating the bank is overwritten at each operation. On the right, a branching architecture is depicted by three banks arranged left to right. A first convolution operation reads from the first bank and writes to a new, second bank. A second convolution reads from all previously written banks and then writes its result to a new, third bank. The process is illustrated by blue arrows from bank 1 and bank 2 that converge on the convolution operation and a red arrow from the convolution operation to the third memory bank. The figure emphasizes that single-branch models reuse and overwrite a single memory bank, while branching models accumulate information by reading from all prior banks and writing to a fresh bank each step.
Figure 5: Example of a memory-bank representation taken from Brock et al. [86]. In a single-branch architecture (LEFT), the network utilizes a singular, large memory bank that it both reads and overwrites at each operation. Branching architectures (RIGHT) read from all previously written banks and write to an empty bank.
Knowledge Distillation
Distillation [87] transfers “teacher” model knowledge into “student” networks (Figure 6). While not yet widely used for SNNaS, SAKD (self-architectural knowledge distillation) [88] shows promise by matching spiking and ANN features under ultra-low latency. Distillation can reduce training costs and improve student generalization, though it adds complexity by requiring balance between teacher–student dynamics.
Diagram depicting the knowledge distillation process. On the left is a large cylinder labeled ‘DATA.’ Two horizontal arrows point from the data to two dashed boxes on the right. The top dashed box is labeled ‘Teacher Model’ and contains a larger neural network, consisting of five fully connected layers of colored nodes. Out of the bottom of this box, a vertical arrow labeled ‘Distill’ points to a rectangle labeled ‘Knowledge.’ From this rectangle, another vertical arrow labeled ‘Transfer’ points down to a second dashed box, containing a smaller neural network of three fully-connected layers. The data cylinder has an arrow into the student box, indicating the student is trained on the same data while also receiving distilled knowledge from the teacher.
Figure 6: Example of knowledge distillation [89].
Hardware Search Space
As network architectures directly affect implementation efficiency (e.g., power, memory), hardware aspects must be considered early in SNNaS. Parameter-based frameworks [90, 91, 84] adjust PEs, buffer sizes, and interfaces, whereas template-based methods [92] reconfigure known hardware design templates. ANCoEF [84] highlights that incorrect neuron-to-PE ratios can waste bits in lookup tables, weight SRAM, and network-on-chip flits. Benmeziane et al. [93] classify hardware into server-grade CPUs/GPUs/FPGAs/ASICs, mobile devices, and tiny devices—each requiring distinct design optimizations.
@@
Memory Usage Refers to the amount of memory required to store model parameters and operations. Mixed-precision quantization (e.g., LitE-SNN [25]); Memory-aware kernel size exploration (e.g., SpikeNAS [12]); Weight pruning and compression (e.g., GAQ-SNN [16])
Latency Time taken to complete inference or training, crucial for real-time applications. Hardware-specific optimization (e.g., HASNAS [22], ANCoEF [84]); Timestep compression in SNNs (e.g., LitE-SNN [25])
Energy Consumption The power required to perform computations, important for low-power devices. Minimizing spike counts (e.g., NeuEvo [18], AutoSNN [7]); Synaptic operation (SynOps)-based evaluation (e.g., Auto-Spikformer [17]); Sparse computation using event-driven architectures.
Hardware Area The physical size of the hardware components required for implementation. Compute-in-memory architectures (e.g., HASNAS [22]); Optimization of neuron and synapse allocations.
Hardware Compatibility Ensures the designed SNN can efficiently operate on specific hardware. Co-exploration of architecture and hardware (e.g., ANAS [77], ANCoEF [84]); Resource-balancing algorithms for neuromorphic platforms (e.g., ANCoEF [84]).
Bandwidth Data transfer capacity between components, critical in parallel architectures. Data partitioning and mapping (e.g., ANCoEF [84]); Efficient routing and scheduling strategies.
@@ III-C Hardware-Aware NAS
Hardware-aware NAS (HW-NAS) seeks to optimize neural network architectures for specific hardware platforms, balancing traditional performance metrics (e.g., accuracy) with constraints like energy, latency, memory usage, and computational requirements [94, 93]. Importantly, temporal and event-driven SNN computations behave differently across hardware implementations [6], so even theoretically optimal SNNs may be impractical if they fail to align with platform-level design choices. As summarized in Table II, HW-NAS ensures that the final architectures exploit hardware features—like event-driven processing or sparse computation [95]—to achieve substantial efficiency gains.
@@
In general, there are two main strategies for HW-NAS: single-objective and multi-objective optimization [93]. Single-objective methods either use a two-stage process—first searching for accuracy, then optimizing for hardware—or incorporate hardware limits as constraints from the onset. Multi-objective frameworks, such as SpikeExplorer [26] and LitE-SNN [25], balance accuracy with hardware metrics (e.g., power, area, and memory). Many SNNaS methods rely on Synaptic Operations (SynOps) to model computational and energy costs [17, 25, 16, 24], while spike-count minimization (e.g.,[7, 23, 18]) reduces power. Quantization[16] and kernel-size optimization [21] further shrink memory footprints. HW-NAS methods like HASNAS [22], NeuEvo [18], GAQ-SNN [16], and Auto-Spikformer [17] show that carefully coupling architecture search with hardware-specific constraints yields high accuracy alongside notable improvements in energy and resource efficiency.
III-D Co-Exploring Neural Architecture Search Space and Hardware Design Space
Co-exploration simultaneously searches both the neural architecture and the hardware design, going beyond standard HW-NAS frameworks by optimizing every dimension of the design jointly [111]. Unlike HW-NAS—which assumes a fixed hardware platform—co-exploration may consider different or new hardware architectures to ensure designs are feasible, efficient, and well-aligned with real-world deployment constraints [94]. This approach typically delivers better performance and lower costs than separately optimizing hardware or SNNs, but at the expense of greater search complexity, substantial GPU hours, and a large carbon footprint [112].
@@ III-D Co-Exploring Neural Architecture Search Space and Hardware Design Space
Table III: Architecture Search Strategies

### v2

@@ III-A Unique Challenges of Applying NAS to SNNs
First, while standard NAS relies on gradient-based optimization, SNNs’ discrete spike functions are inherently non-differentiable. Though surrogate gradient methods exist, they introduce approximation errors and computational overhead that impact large-scale architecture searches [41]. The prevalence of spike-based learning rules like STDP, distinct from gradient-based approaches, further complicates traditional NAS applications.
@@
Third, neuromorphic hardware considerations introduce unique constraints such as limited precision and asynchronous computation [62]. Traditional NAS methods that do not account for these constraints may yield architectures incompatible with neuromorphic platforms.
@@
III-B Search Space Design
The search space for a NAS method is the set of feasible network configurations, typically specified by parameters such as the number and type of layers and the layer hyperparameters (e.g., kernel size, filter sizes, or the number of neurons). Because these dimensions yield billions of potential architectures, the search space can be prohibitively large and often becomes the primary bottleneck for architecture search performance [109]. Moreover, different architectural choices can significantly impact hardware accelerators, affecting power efficiency, memory usage, and throughput [7, 9]. Consequently, it is important to design the search space to balance performance and practical hardware constraints.
The search can be simplified and the search space reduced by incorporating prior knowledge of well-performing architectural patterns. However, doing so may introduce human bias, undermining the discovery of novel solutions. On the other hand, overly constraining the search space might exclude high-performing architectures, resulting in suboptimal performance. Hence, selecting an appropriate search space remains a critical challenge in SNNaS, requiring consideration of both algorithmic and hardware factors.
@@ III-B Search Space Design
Global (Unstructured) Search Spaces
@@
Global search spaces consider the entire network configuration at once, allowing architectural decisions to be made across all layers simultaneously without enforced repetition or modular reuse. They offer tremendous diversity but can be extremely large, which makes efficient exploration challenging. Several SNNaS methods employ global search spaces to maximize architectural flexibility. For instance, MixedSNN [167] searches among different spiking neuron models and thresholds, and Auto-Spikformer [30] combines transformer and SNN parameters. Methods such as Liu and Yi [84], efficient SNN (ESNN) [170], Automatic Spiking Transformer (AutoST) [160], Genetic Algorithm-based Quantization framework for Spiking Neural Networks (GAQ-SNN) [99], and asynchronous neuromorphic hardware architecture search (ANAS) [179] illustrate the complexity of global search—spanning from trillions of spiking configurations to hardware-related parameters like processing elements and buffer sizes.
Sequential Search Spaces
Inspired by early NAS works [187, 12], sequential search spaces encode neural architectures as ordered, layer-by-layer constructions, where each layer is specified as an individual architectural choice (e.g., operation type, neuron model, or kernel size), rather than through repeated modules or hierarchical architectural motifs. Because these spaces do not enforce architectural regularity or parameter sharing, their combinatorial size grows rapidly, substantially increasing search complexity. In spiking neural networks, this challenge is further amplified by temporal dynamics, neuron state variables, and spike-driven computation, rendering purely sequential SNNaS prohibitively expensive in practice. Consequently, sequential search spaces remain largely unexplored in practical SNNaS settings, with most modern methods favoring more structured alternatives that better control computational cost.
Cell-Based Search Spaces
@@
Cell-based approaches define networks at macro and micro levels. Macro-architectures arrange repeated cells, while micro-architectures define operations within each cell. SNASNet (NAS approach for finding better SNN architectures) [68], based on NAS-Bench-201 [44], pioneered cell-based SNN search using directed acyclic graphs (DAGs) of operations (zeroize, skip connection, $1\times 1$ convolution, $3\times 3$ convolution, and $3\times 3$ average pooling). Extensions include spiking spatiotemporal NAS (SSTNAS) [74], SpikeNAS [114], Spikernel [112], NeuroNAS [111], and variations in cell design [100]. SpikeDHS (spike-based differentiable hierarchical search) [29, 31] also employs DAG-based cells. Multi-Attention Differentiable Architecture Search (MA-DARTS) [90] extends the cell-based framework to incorporate multi-dimensional attention blocks that distinguish information across channel, spatial, and temporal dimensions and prioritize spatiotemporal spike patterns. MA-DARTS has minimal parameter overhead and improved accuracy. An example cell-based diagram is shown in Figure 2.
@@
Figure 2: Example of cell with four nodes and five possible operations. The cell is a directed acyclic graph, with each edge representing an operation [44].
Hierarchical Structure Search Spaces
@@
Hierarchical NAS generalizes cell-based approaches and arranges architectures in multi-level motifs, mirroring modular designs like VGGNet [137] or ResNet [54]. Multiscale evolutionary NAS (MSE-NAS) [104] organizes SNNs into microscopic, mesoscopic, and macroscopic levels; AutoSNN [97] and Asynchronous Neuromorphic algorithm/hardware Co-Exploration Framework (ANCoEF) [180] combine macro-level backbones with micro-level spiking blocks. SpikeDHS [29, 31] extends from cell-level to layer-level search, and LitE-SNN [82] adds pruning and mixed-precision quantization. SpikeExplorer [102] employs Bayesian optimization (BO) for multi-objective search. Deep Evolutionary SPIking NEtwork (DESPINE) [5] and Neural circuit Evolution strategy (NeuEvo) [132] further emphasize biologically inspired motifs, excitatory/inhibitory neurons, and feedback connections. Figure 3 illustrates a typical three-level hierarchy.
@@
Hardware Design Search Spaces
As network architectures directly affect implementation efficiency (e.g., power, memory), hardware aspects must be considered early in SNNaS. Parameter-based frameworks [2, 32, 180] adjust processing elements (PEs), buffer sizes, and interfaces, whereas template-based methods [171] reconfigure known hardware design templates. ANCoEF [180] highlights that incorrect neuron-to-PE ratios can waste bits in lookup tables (LUTs), weight SRAM, and network-on-chip flits. Benmeziane et al. [15] classify hardware into server-grade CPUs/GPUs/FPGAs/ASICs, mobile devices, and tiny devices—each requiring distinct design optimizations. When jointly explored, neural and hardware search spaces enable end-to-end optimization of PPA metrics (power, performance, and area) and accuracy at the cost of increased search complexity.
@@
Memory Usage The amount of memory required to store model parameters and operations. Mixed-precision quantization (e.g., LitE-SNN [82], ABC-Arc [182], NeuroNAS [111], ASNPC [157]); Memory-aware kernel size exploration (e.g., SpikeNAS [114], NeuroNAS [111]); Weight pruning and compression (e.g., GAQ-SNN [99], ABC-Arc [182], SPEAR [165], ESL-SNNs [133], ANCoEF [180]); Sparse training from scratch (e.g., ESL-SNNs [133]); Linear regression for SynOps estimation (e.g., SPEAR [165]); DRL target-aware reward (e.g., SPEAR [165]); Neuron hardware unit reuse (e.g., ASNPC [157]).
Latency Time taken to complete inference or training, crucial for real-time applications. Hardware-specific optimization (e.g., NeuroNAS [111], ANCoEF [180], ASNPC [157], SpikeExplorer [102], TNAS-ER [83]); Timestep compression in SNNs (e.g., LitE-SNN [82], DESPINE [5], TNAS-ER [83]); Sparsity-aware event skipping (e.g., TNAS-ER [83]); ANN-assisted time-to-first-spike (TTFS) search (e.g., SpikeExplorer [102]).
Energy Consumption The power required to perform computations, important for low-power devices. Minimizing spike counts (e.g., NeuEvo [132], AutoSNN [97], DESPINE [5], EB-NAS [105], ANCoEF [180], MONAS-ESNN [125]); Synaptic operation (SynOps)-based evaluation (e.g., Auto-Spikformer [30], ESNN [170], SPEAR [165], LitE-SNN [82]); Sparse computation using event-driven architectures (e.g., SpikeX [168]); Linear regression for SynOps estimation (e.g., SPEAR [165]); DRL target-aware reward (e.g., SPEAR [165]); Sparsity-aware weight loading (e.g., SpikeX [168]); ASAHD (e.g., MONAS-ESNN [125]).
Hardware Area The physical size of the hardware components required for implementation. Compute-in-memory architectures (e.g., NeuroNAS [111]); Optimization of neuron and synapse allocations (e.g., ANCoEF [180], ANAS [179], MOO-SNNs [43]); Neuron hardware unit reuse (e.g., ASNPC [157]); Mixed-precision quantization (e.g., GAQ [99]).
Hardware Compatibility Ensures the designed SNN can efficiently operate on specific hardware. Co-exploration of architecture and hardware (e.g., ANAS [179], ANCoEF [180], ASNPC [157], NeuroNAS [111], SpikeExplorer [102]); Resource-balancing algorithms for neuromorphic platforms (e.g., ANAS [179], ANCoEF [180], TNAS-ER [83]); Chip-specific search space (e.g., Liu & Yi [84]).
Bandwidth Data transfer capacity between components, critical in parallel architectures. Data partitioning and mapping (e.g., ANCoEF [180], ANAS [179], SpikeX [168], TNAS-ER [83]); Efficient routing and scheduling strategies (e.g., ANAS [179], SpikeX [168]); Compute-in-memory architectures (e.g., NeuroNAS [111]); Sparsity-aware weight loading (e.g., SpikeX [168]); Activity-based data fetching prioritization (e.g., SpikeX [168]).
Neuron-Level Search Spaces
In SNNaS, neuron-level search spaces optimize spiking neurons to suppress redundant spiking activity, ensuring evolved architectures are more accurate and energy-efficient. While some SNNaS works select from different neuron model types tailored to their task [167, 178], others optimize neuron parameters that might include leakage, decay, spiking frequencies, adaptive thresholds, time steps, reset mechanisms, and even surrogate gradient shapes to better capture temporal dynamics [5, 170, 29, 31, 30, 128, 25, 186, 104, 157].
@@ III-C Hardware-Aware NAS
Hardware-aware NAS (HW-NAS) seeks to optimize neural network architectures for specific fixed target hardware instantiations (e.g., GPUs, FPGAs, mixed-signal neuromorphic hardware, asynchronous platforms, and emerging hybrid architectures), balancing traditional performance metrics (e.g., accuracy) with hardware-related constraints like energy, latency, memory usage, and computational requirements, without altering the underlying neural architecture search space [184, 15]. Importantly, temporal and event-driven SNN computations behave differently across hardware implementations [108], so even theoretically optimal SNNs may be impractical if they fail to align with platform-level design choices. As summarized in Table II, HW-NAS ensures that the final architectures exploit hardware features—like event-driven processing or sparse computation [117]—to achieve substantial efficiency gains.
@@
In general, there are two main strategies for HW-NAS: single-objective and multi-objective optimization [15]. Single-objective methods either use a two-stage process—first searching for accuracy, then optimizing for hardware—or incorporate hardware limits as constraints from the onset. Multi-objective frameworks, such as SpikeExplorer [102] and LitE-SNN [82], balance accuracy with hardware metrics (e.g., power, area, and memory). Many SNNaS methods rely on Synaptic Operations (SynOps) to model computational and energy costs [30, 82, 99, 170], while spike-count minimization (e.g.,[97, 105, 132]) reduces power. Quantization [99] and kernel-size optimization [112] further shrink memory footprints. HW-NAS methods like NeuroNAS [111], NeuEvo [132], GAQ-SNN [99], and Auto-Spikformer [30] show that carefully coupling architecture search with hardware-specific constraints yields high accuracy alongside notable improvements in energy and resource efficiency.
III-D Bridging the Hardware-Software Co-Design Gap in SNNaS
III-D1 Co-Exploring the Neural Architecture Search Space and Hardware Design Space
Co-exploration simultaneously searches neural architectures and hardware design parameters, going beyond standard HW-NAS frameworks by optimizing every dimension of the design jointly [64]. Unlike HW-NAS, which assumes a fixed hardware platform, co-exploration may consider different or new hardware architectures to ensure designs are feasible, efficient, and well-aligned with real-world deployment constraints [184]. This approach typically delivers better performance and lower costs than separately optimizing hardware or SNNs, but at the expense of greater search complexity, substantial GPU hours, and a large carbon footprint [141].
@@ III-D1 Co-Exploring the Neural Architecture Search Space and Hardware Design Space
Studies show neuromorphic hardware-software co-design achieves breakthrough performance in narrow, well-defined domains where end-to-end optimization is possible [180, 168, 157, 5], but fails systematically when architectural misalignment occurs or communication bottlenecks are present [26, 14, 108, 129].
III-D2 The Hardware/Software Boundary
Perhaps the most critical frontier in SNNaS is the "hardware/software boundary," the interface where mathematical abstractions must reconcile with the physical realities of neuromorphic substrates [42, 127, 111]. Standard NAS approaches typically treat hardware as a fixed penalty term in a fitness function, but this misses a crucial insight: synchronous (clock-driven) and asynchronous (event-driven) hardware require fundamentally different architectural paradigms [42]. Performance is often further constrained by communication overhead between software simulators and physical chips [127].
III-D3 Asynchronous and Event-Driven Hardware Constraints
Current asynchronous, event-driven hardware platforms like Intel’s Loihi and Synsense’s Speck achieve their power efficiency by eliminating the global clock, but this design choice imposes stringent architectural constraints [42, 84]. Without clock synchronization, spike arrival times become imprecise, making certain operations problematic for existing asynchronous neuromorphic chips. Spiking matrix multiplication (as used in Spiking Transformers) and max-pooling layers are particularly vulnerable—even small timing delays on contemporary asynchronous neuromorphic platforms can produce significant errors in attention maps [42]. Hardware-specific limitations compound these challenges; Loihi, for instance, lacks support for backward and shortcut connections [84].
The Spiking Token Mixer (STMixer) [42] navigates these constraints by restricting its search to asynchronous-compatible operations—convolutions, fully connected layers, and residual paths—while deliberately avoiding operations that demand precise spike synchronization. ANAS [179] tackles the problem from the hardware side, using CanMore to search across both numerical parameters (PE count, buffer size) and structural choices (NoC topologies, routing algorithms), optimizing how Address-Event Representation protocols and axon-sharing rules are implemented to match SNN sparsity patterns.
Table III: Spiking Neural Network Architecture Search Strategies